# LLMs Cannot Reliably Judge (Yet?): A Comprehensive Assessment on the Robustness of LLM-as-a-Judge

### v1

@@ 4.8 Content Quality Evaluator
where $x$ denotes the target response (either $r$ or $r_{m}$), and $r_{\text{ref}}$ is the corresponding reference response. We design specialized evaluators for different task categories, i.e., text-focused, code-oriented evaluation tasks, to ensure appropriate quality assessment.
Evaluator for Text-focused Tasks $Q_{\text{text}}$. For text-focused evaluation tasks such as summarization and translation, we assess the fluency, coherence, and relevance of the response using BLEURT [54], a state-of-the-art BERT-based model fine-tuned on human-labeled data. BLEURT computes semantic similarity between two responses, returning a score in the range (0,1). The reference responses $r_{\text{ref}}$ are drawn from benchmark datasets corresponding to each task.
Evaluator for Code-oriented Tasks $Q_{\text{code}}$. For code-oriented tasks, we introduce an evaluator to assess functional correctness, syntactic validity, and semantic accuracy. We adopt CodeBLEU  [55], a metric designed for programming languages. CodeBLEU extends the traditional BLEU score by incorporating code-specific properties, such as syntax structure, data-flow consistency, and token-level similarity. The final CodeBLEU score also lies in the range (0,1).
Evaluator for Knowledge-intensive Tasks $Q_{\text{know}}$. For knowledge-intensive tasks, such as logical reasoning and mathematical problem solving, semantic similarity measures fail to accurately reflect content quality. In these cases, the correspondence between the model response and the reference often involves abstract reasoning rather than surface-level similarity. As such, we disable the content quality evaluator for these tasks and assign a fixed score of $S_{e}=0$.
4.9 Evaluation Metrics
@@
    $\text{SDR}=\frac{1}{N}\sum_{i=1}^{n}(s_{t}^{(i)}-\hat{s}_{t}^{(i)})\times 0.1$ (9)
    where $N$ denotes the size of the test set. $\hat{s}_{t}^{(i)}$ is the original score for instance $i$, and $s_{t}^{(i)}$ is the score after attack or defense. The factor of 0.1 serves to normalize the $\hat{s}_{t}^{(i)}$ and $s_{t}^{(i)}$ value to the range (0, 1).
  • Improved Score Difference Rate (iSDR): Since adversarial modifications or defenses may enhance the quality of responses (leading to unintended score improvements) as a side effect, we propose iSDR to isolate the impact of manipulation from genuine content improvement. iSDR compensates for content quality changes by subtracting the corresponding content quality score change. Formally:
@@
    $=\frac{1}{N}\sum_{i=1}^{n}\left((s_{t}^{(i)}-\hat{s}_{t}^{(i)})\times 0.1-(s_{e}^{(i)}-\hat{s}_{e}^{(i)})\right)$
@@
  • Attack Success Rate (ASR): ASR measures the proportion of successful adversarial attacks, defined differently for pointwise and pairwise evaluation protocols. Pointwise ASR (ASR) is defined as the proportion of test cases with positive iSDR, representing a manipulated judgment:
@@
    Pairwise ASR (P-ASR) considers both of the candidate orders and measures how often the judge fails to select the reference response $r_{\text{ref}}$ as the better choice.
[…]
@@
  • Accuracy, Recall, Precision, and F1: These metrics are used to evaluate the performance of detection-based defense techniques, which aim to distinguish between adversarial and benign responses. Specifically, the defense guard attempts to correctly identify adversarial responses (TP), preserve benign responses (TN), and avoid misclassifying a benign response as adversarial (FP) or failing to detect an adversarial response (FN). Based on these quantities, we compute Accuracy (ACC), Recall, Precision (Pre), and F1 Score.
TABLE III: Evaluation of adversarial attack for the responses generated by openchat-3.5 on translation tasks. Higher values indicate superior attack effectiveness for all metrics.
@@ 5 Experiments
In this section, we leverage RobustJudge to systematically compare the effectiveness of different adversarial attack methods against the LLMs listed in Table XI [^1] We accessed all models via official APIs or hosted endpoints provided by Together.ai, HuggingFace, and OpenAI.. Our evaluation aims to answer the following research questions:
  • RQ1: What impact do different adversarial attacks and defense methods have on the LLM-based judges?
  • RQ2: How do the prompt templates and model choices affect the robustness of LLM-based judges?
  • RQ3: What vulnerabilities persist in black-box real-world deployments of LLM-based judges, as revealed by our empirical evaluations?
5.1 Attack and Defense Effectiveness (RQ1)
5.1.1 Attack Performance
We conduct all main experiments on Openchat-3.5 and present results on other judge models (e.g., Qwen-2.5) in Table XIII (appendix).
Impact of Advarsarial Attacks. We evaluate the effectiveness of various adversarial attack methods, and report their ASR, SDR, iSDR and P-ASR in Table III.
@@ 5.1.1 Attack Performance
Heuristic-based attacks generally exhibit strong and stable effectiveness across all evaluated models. For instance, most heuristic attacks (H1–H7) consistently achieve high ASRs, frequently exceeding 80%. Notably, the Combined Attack (H6) achieves 100% ASR on Openchat-3.5, Qwen-2.5, and Mistral-7B, and above 90% on Llama-3.1-8B and Llama-3.3-70B. This trend is also reflected in SDR and iSDR, where H6 yields the highest SDR values, e.g., 0.5510 (Openchat-3.5), 0.5560 (Qwen-2.5), 0.5652 (Mistral-7B), 0.5985 (Llama-3.1-8B), and 0.5160 (Llama-3.3-70B). Additionally, H6 exhibits high impact across different underlying judge model. We attribute the effectiveness of H6 to its multi-faceted attack strategy, which integrates multiple manipulations, such as escape characters, context manipulation, and adversarial completions, into a single composite prompt. The attack succeeds if any one of its constituent mechanisms is effective, allowing its effects to accumulate. These results suggest that combining compatible attack techniques can lead to amplified adversarial effectiveness, highlighting the importance of defense mechanisms that address multi-facet threat scenarios.
@@
Optimization-based attacks are also effective against LLM-as-a-Judge systems, particularly methods O1 (AdvEval), O2 (PAIR), and O3 (TAP). For instance, O1 achieves 86.67% ASR on Qwen-2.5, 90% ASR on Mistral-7B, and 93.33% on Llama-3.3-70B. Similarly, O2 attains 100% ASR on Qwen-2.5 and 90% on Llama-3.1-8B, while O3 also yields consistently high ASR scores across multiple models. In addition to ASR, we observe a similar trend on the other metrics such as SDR and iSDR, which further confirms the effects of these attacks. In contrast, other optimization-based approaches (O5–O7, e.g., GCG, AutoDAN, Greedy) result in lower ASRs (typically below 50%) and exhibit limited or even negative SDR/iSDR improvement. These results demonstrate their relatively weak optimization effectiveness and poor cross-model transferability. These results also reflect our experimental constraints. Given the high computational cost of optimization-based attacks, we restrict the number of iterations for each run. Although this may affect the final performance, it ensures practical feasibility and fair comparison.
$\bullet$ Finding 1: Existing LLM-as-a-Judge systems are still highly vulnerable to adversarial attacks, particularly methods Fake Reasoning (H5), Combined Attack (H6), Empty (H7) and AdvEval (O1), PAIR (O2), TAP (O3).
TABLE IV: Evaluation results across multiple tasks (T1–T8) and models under three representative attack methods: Combined Attack (H6), Fake Reasoning (H5), and Fake Completion (H4).
Type Task Judge Model Attack Methods
Combined Attack (H6) Fake Reasoning (H5) Fake Completion (H4)
ASR (%) SDR iSDR P-ASR (%) ASR (%) SDR iSDR P-ASR (%) ASR (%) SDR iSDR P-ASR (%)
Text T1 openchat-3.5 100.00 0.551 0.290 28.33 90.00 0.442 0.181 40.00 40.00 0.221 -0.040 10.00
Qwen2.5-7B 100.00 0.556 0.476 60.00 90.00 0.356 0.276 28.33 36.67 0.019 -0.061 21.67
Mistral-7B 100.00 0.565 0.491 63.33 89.66 0.446 0.372 48.33 76.67 0.323 0.248 66.67
LLama-3.1-8B 90.00 0.599 0.601 71.67 60.00 0.146 0.149 46.67 60.00 0.051 0.054 46.67
T2 openchat-3.5 90.00 0.491 0.258 45.00 85.00 0.438 0.204 90.00 85.00 0.466 0.232 67.50
Qwen2.5-7B 95.00 0.541 0.346 55.00 95.00 0.480 0.285 87.50 100.00 0.457 0.262 52.50
Mistral-7B 90.00 0.491 0.207 97.50 90.00 0.509 0.224 90.00 95.00 0.519 0.234 100.00
LLama-3.1-8B 80.00 0.370 0.150 75.00 80.00 0.365 0.145 65.00 90.00 0.373 0.154 35.00
Code T3 openchat-3.5 83.33 0.593 0.100 35.00 63.33 0.508 0.016 31.67 73.33 0.625 0.133 41.67
Qwen2.5-7B 93.33 0.648 0.288 0.00 70.00 0.438 0.078 56.67 56.67 0.411 0.051 61.67
Mistral-7B 83.33 0.734 0.163 98.33 83.33 0.740 0.169 80.00 83.33 0.700 0.129 96.67
LLama-3.1-8B 86.67 0.641 0.136 55.00 66.67 0.507 0.001 33.33 63.33 0.531 0.025 51.67
T4 openchat-3.5 85.00 0.716 0.118 10.00 65.00 0.698 0.100 50.00 80.00 0.723 0.125 60.00
Qwen2.5-7B 95.00 0.742 0.143 22.50 70.00 0.659 0.061 95.00 80.00 0.703 0.105 97.50
Mistral-7B 100.00 0.742 0.141 100.0 100.00 0.768 0.167 87.50 100.00 0.761 0.160 90.00
LLama-3.1-8B 95.00 0.742 0.160 87.50 80.00 0.653 0.072 60.00 75.00 0.653 0.071 77.50
T5 openchat-3.5 100.00 0.593 0.199 45.00 90.00 0.552 0.158 92.50 95.00 0.608 0.214 70.00
Qwen2.5-7B 35.00 0.602 0.008 57.50 45.00 0.613 0.018 57.50 45.00 0.581 -0.013 62.50
Mistral-7B 100.00 0.581 0.080 100.0 95.00 0.551 0.051 97.50 95.00 0.618 0.117 97.50
LLama-3.1-8B 100.00 0.581 0.264 82.50 95.00 0.485 0.168 92.50 95.00 0.553 0.236 60.00
Knowledge T6 openchat-3.5 70.00 0.255 0.229 29.46 75.00 0.242 0.216 27.68 85.00 0.264 0.238 35.00
Qwen2.5-7B 95.00 0.546 0.417 1.790 95.00 0.320 0.192 14.29 50.00 0.134 0.005 20.00
Mistral-7B 90.00 0.950 0.220 77.50 95.00 0.938 0.208 57.50 90.00 0.865 0.135 60.00
LLama-3.1-8B 100.00 0.998 0.306 35.00 75.00 0.824 0.133 7.500 65.00 0.772 0.081 17.50
T7 openchat-3.5 65.00 0.359 0.348 23.47 60.00 0.356 0.345 20.41 40.00 0.138 0.126 17.50
Qwen2.5-7B 95.00 0.369 0.342 1.020 45.00 0.170 0.143 32.65 60.00 0.097 0.069 10.00
Mistral-7B 85.00 0.900 0.223 85.00 70.00 0.768 0.091 55.00 70.00 0.760 0.083 57.50
LLama-3.1-8B 85.00 0.850 0.163 62.50 65.00 0.765 0.078 10.00 60.00 0.756 0.069 30.00
T8 openchat-3.5 85.71 0.445 0.437 17.86 89.21 0.413 0.406 22.40 50.00 0.120 0.112 23.21
Qwen2.5-7B 35.71 0.183 0.178 13.33 10.71 0.048 0.043 38.33 $0$0.004 -0.008 35.71
Mistral-7B 53.57 0.718 0.252 73.21 50.00 0.650 0.184 44.64 39.29 0.455 -0.011 64.29
LLama-3.1-8B 85.71 0.921 0.301 41.07 67.86 0.853 0.233 21.43 71.43 0.856 0.236 41.07
@@
Attacks on different Tasks Categories. We conduct evaluation on a variety of tasks, e.g., text-focused, code-oriented and knowledge-intensive tasks, to examine how vulnerabilities to adversarial attacks varies by task categories. The results are presented in Table IV. Additional results on other tasks are provided in Tables XII, XIV, XV, XVI, XVII, XVIII, and XIX in the appendix.
@@
We observe substantial differences in ASR across these tasks. Text-focused tasks (T1–T2), such as machine translation and summarization, exhibit the highest vulnerability. For example, under the Combined Attack, all judge models demonstrate ASRs exceeding 90% on machine translation (T1). In contrast, knowledge-intensive tasks, such as mathematical problem solving (T7) and knowledge recall (T8) show significantly greater robustness. For instance, Qwen2.5-7B achieves an ASR of only 0% on T8 under the Fake Completion attack, and just 45% ASR on T7 under Fake Reasoning attack (H5). This observation highlights the need for task-aware defense mechanisms, especially when deploying LLM-as-a-Judge systems on text-focused tasks.
@@
We believe these discrepancies stem from fundamental differences in task complexity and evaluation depth. Text-focused tasks often rely on surface-level features, such as fluency or coherence, which adversarial attacks can easily manipulate. As a result, LLM judges may be misled by outputs that appear syntactically sound but are semantically or factually incorrect. In contrast, knowledge-intensive tasks require deeper semantic understanding, logical reasoning, or domain-specific knowledge accuracy. These tasks are less vulnerable to superficial prompt modification. This analysis highlights the importance of task-aware defense strategies, particularly when deploying LLM-as-a-Judge systems on text-focused tasks.
@@
$\bullet$ Finding 2: Tasks that emphasize surface-level features (e.g., text-forcused task) exhibit greater vulnerability than knowledge-intensive tasks, which require deeper semantic understanding and reasoning.
@@
Comparison of Judge Protocals. In this section, we compare two evaluation protocols used on LLM-as-a-Judge systems: pointwise scoring and pairwise comparison. The corresponding results are shown in Table III and Figure 5.
@@
Our findings show that while many adversarial strategies achieve high ASR under pointwise scoring, their P-ASR drops considerably in the pairwise comparison setting. For instance, Combined Attack (H6) achieves a perfect ASR of 100% against Openchat-3.5, yet its P-ASR is only 28.33%. Similar trends are observed for Qwen2.5-7B-Instruct (ASR: 100%, P-ASR: 60.00%) and Mistral-7B (ASR: 100%, P-ASR: 63.33%). This discrepancy can be attributed to the nature of the pairwise protocol, which provides the judge model with an additional reference response for comparison. This allows the model to contrast the manipulated response with a stable reference, thereby reducing the attack’s impact.
@@
Despite the absolute differences in ASR and P-ASR, the relative effectiveness of attacks remains consistent for both protocols. Combined Attack (H6) and Fake Reasoning (H5) consistently rank among the most effective attacks in both ASR and P-ASR against all judge models, as shown in Table III. Conversely, optimization-based attack GCG (O5) and Auto-DAN (O6) exhibit limited impact on Openchat-3.5 (e.g., 0% and 6.67% ASR, respectively) under pointwise protocol, and correspondingly low P-ASRs (e.g., P-ASR <10% in most cases) under the pairwise protocol.
@@
TABLE V: Evaluation of prevention-based defense methods (D1-D4) against adversarial attacks.
@@
D1 ASR(%) 0.000 0.000 0.000 0.000 3.330 100.00 90.00 3.330 3.330 3.330 3.330 23.33 6.670 6.670 6.670 16.67
D2 ASR(%) 73.33 80.00 80.00 53.33 90.00 96.67 93.33 40.00 76.67 76.67 80.00 33.33 6.670 10.00 20.00 60.67
D3 ASR(%) 76.67 86.67 86.67 46.67 96.67 100.0 86.67 50.00 83.33 83.33 90.00 33.33 0.000 6.670 23.33 63.33
D4 ASR(%) 80.00 80.00 80.00 46.67 93.33 100.0 90.00 43.33 80.00 76.67 73.33 26.67 13.33 10.00 26.67 61.33
TABLE VI: Evaluation of detection-based defense methods (D5–D7) against adversarial attacks.
Defense Metric Attack Methods Average
H1 H2 H3 H4 H5 H6 H7 H8 O1 O2 O3 O4 O5 O6 O7
D5 ACC(%) 55.00 53.33 51.67 50.00 48.33 50.00 71.67 23.64 46.67 31.67 28.33 51.67 51.67 55.00 56.67 48.35
@@
5.1.2 Defense Performance
We evaluate 7 representative defenses methods from both prevention-based and detection-based paradigms. These include: retokenization[17] (D1), delimiter insertion (D2), sandwich prompting (D3), instruction augmentation (D4), perplexity filtering (PPL) (D5), windowed PPL (WinPPL) (D6), and a naive LLM-based detector (D7). Since prevention-based and detection-based methods are based on different mechanism and evaluated on different metrics, we report their results separately in Table V and Table VI, respectively.
Prevention-based Defenses (D1–D4). Overall, prevention-based methods demonstrate moderate effectiveness across most attacks. However, they remain largely ineffective against stronger adversarial attacks, such as Combined Attack (H6) and AdvEval (H7). Among these methods, Retokenization (D1) yields the best average defense performance, achieving a reduced ASR of 16.67% and an iSDR of -0.394, significantly outperforming the other methods (D2–D4), which exhibit ASRs exceeding 60% and positive iSDR values above 0.03. Despite this relative success, prior findings [17] indicate that Retokenization (D1) can degrade model performance on benign inputs. This degradation arises from altered tokenization patterns, potentially affecting how the model interprets the content and leading to inaccurate evaluations. Thus, while Retokenization offers strong adversarial robustness, it introduces a trade-off between defense effectiveness and benign input performance.
Detection-based Defenses (D5–D7) Among detection-based methods, perplexity-based filters (PPL and WinPPL) generally underperform. They achieve relatively low accuracy (48.35% and 51.22%), recall (53.11% and 62.44%), precision (45.01% and 48.66%), and F1 scores (0.518 and 0.545), indicating limited capability in distinguishing adversarial inputs. In contrast, the naive LLM-based detector achieves the highest detection performance across all metrics, with an accuracy (65.67%), recall (64.71%), precision (65.49%) and F1 score (62.68%), outperforming all other prevention-based defenses, particularly against complex heuristic-based manipulations. Its effectiveness stems from leveraging a powerful LLM to identify subtle adversarial patterns in responses. However, this approach incurs significant computational overhead, as it requires querying the LLM detector for each input response.

### v2

@@ IV-D2 Content Quality Evaluator
where $x$ denotes the target response (either clean response $r$ or modified response $r_{m}$), and $r_{\text{ref}}$ is the corresponding reference response from benchmark datasets. We design task-specific evaluators to ensure appropriate quality assessment across different domains.
Evaluator for Text-focused Tasks. For text-focused evaluation tasks such as summarization and translation, we assess the semantic similarity and coherence of responses using BLEURT bleurt , a BERT-based metric fine-tuned on human judgments. BLEURT computes semantic alignment between responses and returns scores in [0,1]. This metric effectively captures content-level deviations while remaining efficient for large-scale evaluation. The reference responses $r_{\text{ref}}$ are drawn from benchmark datasets corresponding to each task.
Evaluator for Code-oriented Tasks. For code-oriented tasks including code generation, translation, and summarization, we employ CodeBLEU codebleu , a metric specifically designed for programming languages. CodeBLEU extends traditional BLEU scores by incorporating code-specific properties such as abstract syntax tree matching, data-flow consistency, and token-level similarity. This enables reliable detection of both syntactic and semantic changes in code. The final CodeBLEU score lies in [0,1].
Evaluator for Knowledge-intensive Tasks. For knowledge-intensive tasks, such as mathematical problem-solving and logical reasoning, surface-level similarity metrics often fail to capture correctness. These tasks require symbolic computation and abstract reasoning, where token-level alignment does not reflect semantic validity.
For mathematical tasks, we compute exact match accuracy by comparing the final answer against ground-truth solutions, yielding $S_{e}\in\{0,1\}$. For logical reasoning tasks, we employ rule-based consistency checkers that verify logical validity and structural coherence, producing binary correctness scores. In cases where gold-standard references are unavailable or automated evaluation remains unreliable, we conservatively set $S_{e}=0$, treating content as unchanged.
IV-D3 Evaluation Metrics
$s=M(P_{1},r)$ (6)
Pairwise Evaluation. This protocol compares two candidate responses and determines which one is preferred in the given context. The model outputs a preference for either response $r_{a}$ or $r_{b}$, depending on the relative quality of the two responses. Formally, this can be expressed as:
$p=M(P_{2},r_{a},r_{b})$ (7)
where $P_{2}$ is the pairwise prompt template, and $r_{a}$ and $r_{b}$ are the two responses under evaluation, $p\in\{r_{a},r_{b}\}$. Prior research on LLM-as-a-Judge systems judgeFirst has demonstrated that the input order of candidate responses can influence the evaluation outcome. To account for this bias, evaluations are typically conducted using both input orders, with the results defined as:
$p_{+}=M(P_{2},r_{a},r_{b}),\quad p_{-}=M(P_{2},r_{b},r_{a})$ (8)
@@ IV-D3 Evaluation Metrics
    $\text{SDR}=\frac{\zeta}{N}\sum_{i=1}^{n}(s_{t}^{(i)}-\hat{s}_{t}^{(i)})$ (9)
    Where $N$ denotes the size of the test set. $\hat{s}_{t}^{(i)}$ is the original score for instance $i$, and $s_{t}^{(i)}$ is the score after attack or defense. We set $\zeta=0.1$ to ensure that both $\hat{s}_{t}^{(i)}$ and $s_{t}^{(i)}$ are normalized to lie within the range (0, 1).
  • Improved Score Difference Rate (iSDR): Since adversarial modifications or defenses may enhance the quality of responses (leading to unintended score improvements) as a side effect, we propose iSDR to isolate the impact of manipulation from genuine content improvement. iSDR compensates for changes in content quality by subtracting the corresponding change in content quality score. Formally:
@@
    $=\frac{1}{N}\sum_{i=1}^{n}\left(\zeta(s_{t}^{(i)}-\hat{s}_{t}^{(i)})-(s_{e}^{(i)}-\hat{s}_{e}^{(i)})\right)$
@@
  • Attack Success Rate (ASR): ASR measures the proportion of successful adversarial attacks, defined differently for pointwise and pairwise evaluation protocols. Pointwise ASR is defined as the proportion of test cases with positive iSDR, representing a manipulated judgment:
@@
    Pairwise ASR considers both of the candidate orders and measures how often the judge fails to select the reference response $r_{\text{ref}}$ as the better choice.
    $\text{P-ASR}=\frac{1}{2N}\sum_{i=1}^{n}\left(\mathbbm{1}(p_{+}\neq r_{\text{ref}})+\mathbbm{1}(p_{-}\neq r_{\text{ref}})\right)$ (12)
@@
TABLE III: Evaluation of adversarial attacks on translation responses generated by OpenChat-3.5, judged by five different LLM-based evaluators. Columns H1–H8 correspond to handcrafted attacks: H1 Naive Attack naive_2 , H2 Escape Characters injection , H3 Context Ignoring context_ignore_1 ; context_ignore_2 , H4 Fake Completion fake_completion , H5 Fake Reasoning judgeDeceiver , H6 Combined Attack prompt_bench , H7 Empty Prompt, H8 Long Suffix. Columns O1–O7 denote optimization-based or automated attacks: O1 AdvEval adv , O2 PAIR pair , O3 TAP tap , O4 Cheating cheating , O5 GCG gcg , O6 AutoDAN autodan , O7 Greedy Attack op2 . Metrics include attack success rate (ASR↑), semantic degradation rate (SDR↑), and preference-aligned success rate (P-ASR↑). Higher values indicate stronger attack effectiveness.
@@ V Experiment
In this section, we leverage RobustJudge to systematically compare the effectiveness of different adversarial attack methods against the LLMs listed in Table IV [^4] We accessed all models via official APIs or hosted endpoints provided by Together.ai, HuggingFace, and OpenAI.. Our evaluation aims to answer the following research questions.
TABLE IV: Characteristics of the studied LLM judges. The models are grouped by type: closed-source, base open-source, fine-tuned open-source, and reasoning-specialized.
Judge Model Backbone Params Fine- tuned Reasoning Open Source
Closed-Source Models
GPT-4o GPT-4o -
PAI-Judge Qwen -
Base Open-Source Models
Llama-3.1 Llama 8B
Llama-3 Llama 70B
Openchat-3.5 Mistral 7B
Qwen-2.5 Qwen-2.5 7B
Mistral Mistral 7B
Fine-tuned Open-Source Models
JudgeLM Llama 7B,13B
PandaLM Llama-3.1 8B
Auto-J Llama-3.1 8B
Prometheus 2 Mistral 7B
Reasoning-Specialized Model
DS-R1 DS-V3 -
V-A Attack and Defense Effectiveness (RQ1)
V-A1 Attack Performance
We evaluate the attack performance from several aspects.
@@ V-A1 Attack Performance
Impact of Adversarial Attacks. As shown in Table III, heuristic-based attacks generally exhibit strong and stable effectiveness across all evaluated models. For instance, most heuristic attacks (H1–H7) consistently achieve high ASRs, frequently exceeding 8080%. Notably, the Combined Attack (H6) achieves 100% ASR on Openchat-3.5, Qwen-2.5-7B, and Mistral-7B, and above 9090% on Llama-3.1-8B and Llama-3.3-70B. This trend is also reflected in SDR and iSDR, where H6 yields the highest SDR values, e.g., $0.551$ (Openchat-3.5), $0.556$ (Qwen-2.5-7B), $0.565$ (Mistral-7B), $0.599$ (Llama-3.1-8B), and $0.516$ (Llama-3.3-70B). Additionally, H6 exhibits high impact across different underlying judge models. We attribute the effectiveness of H6 to its multi-faceted attack strategy, which integrates multiple manipulations, such as escape characters, context manipulation, and adversarial completions, into a single composite prompt. The attack succeeds if any one of its constituent mechanisms is effective, allowing its effects to accumulate. These results suggest that combining compatible attack techniques can lead to amplified adversarial effectiveness, highlighting the importance of defense mechanisms that address multi-facet threat scenarios.
@@
Optimization-based attacks are also effective against LLM-as-a-Judge systems, particularly methods O1 (AdvEval), O2 (PAIR), and O3 (TAP). For instance, O1 achieves 83.3383.33% ASR on Qwen-2.5-7B, 9090% ASR on Mistral-7B, and 93.3393.33% on Llama-3.3-70B. Similarly, O2 attains 100% ASR on Qwen-2.5-7B and 9090% on Llama-3.1-8B, while O3 also yields consistently high ASR scores across multiple models. In addition to ASR, we observe a similar trend on the other metrics, such as SDR and iSDR, which further confirms the effects of these attacks. In contrast, other optimization-based approaches (O5–O7, e.g., GCG, AutoDAN, Greedy) result in lower ASRs (typically below 50%) and exhibit limited or even negative SDR/iSDR improvement. These results demonstrate their relatively weak optimization effectiveness and poor cross-model transferability. These results also reflect the limitations of our experimental constraints. Given the high computational cost of optimization-based attacks, we restrict the number of iterations for each run. Although this may affect the final performance, it ensures practical feasibility and fair comparison.
$\bullet$ Finding 1: Existing LLM-as-a-Judge systems exhibit systemic vulnerability to adversarial prompts, with certain composite strategies (e.g., Combined Attack) consistently emerging as the most disruptive across models. Elevated SDR and ISDR values further confirm that such attacks substantially distort the semantic integrity of the judged input.
Figure 3: Evaluation results across multiple tasks (T1–T8). The tasks include: T1 (Text Translation), T2 (Text Summarization), T3 (Code Translation), T4 (Code Generation), T5 (Code Summarization), T6 (Logical Reasoning), T7 (Mathematics), and T8 (Knowledge Recall).
@@
Attacks on Different Tasks Categories. We conduct evaluations on a variety of tasks, including text-focused, code-oriented, and knowledge-intensive tasks, to examine how vulnerabilities to adversarial attacks vary by task category. The results are presented in Figure 3.
@@
We observe substantial differences in ASR across these tasks. Text-focused tasks (T1–T2), such as machine translation and summarization, exhibit the highest vulnerability. For example, under the Combined Attack, all judge models demonstrate ASRs exceeding 90% on machine translation (T1). In contrast, knowledge-intensive tasks, such as mathematical problem-solving (T7) and knowledge recall (T8), show significantly greater robustness. For instance, Qwen2.5-7B achieves an ASR of only 0% on T8 under the Fake Completion attack, and 45% ASR on T7 under Fake Reasoning attack (H5). This observation underscores the necessity for task-aware defense mechanisms, particularly when deploying LLM-as-a-Judge systems on text-focused tasks.
@@
We believe these discrepancies stem from fundamental differences in task complexity and the depth of evaluation. Text-focused tasks often rely on surface-level features, such as fluency or coherence, which adversarial attacks can easily manipulate. As a result, LLM judges may be misled by outputs that appear syntactically sound but are semantically or factually incorrect. In contrast, knowledge-intensive tasks require deeper semantic understanding, logical reasoning, or domain-specific knowledge accuracy. These tasks are less vulnerable to superficial prompt modification. This analysis highlights the importance of task-aware defense strategies, particularly when deploying LLM-as-a-Judge systems on text-focused tasks.
@@
$\bullet$ Finding 2: Tasks that emphasize surface-level features (e.g., text-focused tasks) exhibit greater vulnerability than knowledge-intensive tasks, which require deeper semantic understanding and reasoning.
@@
Comparison of Judge Protocols. In this section, we compare two evaluation protocols used on LLM-as-a-Judge systems: pointwise scoring and pairwise comparison. The corresponding results are shown in Table III.
@@
Our findings show that while many adversarial strategies achieve high ASR under pointwise scoring, their P-ASR drops considerably in the pairwise comparison setting. For instance, Combined Attack (H6) achieves a perfect ASR of 100% against Openchat-3.5, yet its P-ASR is only 28.33%. Similar trends are observed for Qwen2.5-7B-Instruct (ASR: 100%, P-ASR: 60.00%) and Mistral-7B (ASR: 100%, P-ASR: 63.33%). This discrepancy can be attributed to the nature of the pairwise protocol, which provides the judge model with an additional reference response for comparison.
@@
Despite the absolute differences in ASR and P-ASR, the relative effectiveness of attacks remains consistent for both protocols. Combined Attack (H6) and Fake Reasoning (H5) consistently rank among the most effective attacks in both ASR and P-ASR against all judge models, as shown in Table III. Conversely, optimization-based attack GCG (O5) and Auto-DAN (O6) exhibit limited impact on Openchat-3.5 (e.g., 0% and 6.67% ASR, respectively) under the pointwise protocol, and correspondingly low P-ASRs (e.g., P-ASR <10% in most cases) under the pairwise protocol.
@@
TABLE V: The results of three detection-based defense methods (PPL, Windows PPL, and Naive LLM) evaluated against handcrafted attacks (H1–H8) and optimization-based attacks (O1–O7). The reported metrics include accuracy (ACC, %), recall (%), precision (%), and F1 score.
@@
PPL ACC(%) 55.00 53.33 51.67 50.00 48.33 50.00 71.67 23.64 46.67 31.67 28.33 51.67 51.67 55.00 56.67 48.35
@@
TABLE VI: The results of four prevention-based defense methods (Retokenizer, Delimiters, Sandwich Prevention, and Instructional Prevention) evaluated against handcrafted attacks (H1–H8) and optimization-based attacks (O1–O7). The reported metrics include attack success rate (ASR, %), score difference rate (SDR), and improved score difference rate (iSDR).
Defense Metric Attack Methods Average
H1 H2 H3 H4 H5 H6 H7 H8 O1 O2 O3 O4 O5 O6 O7
Retokenizer ASR(%) 0.000 0.000 0.000 0.000 3.330 100.00 90.00 3.330 3.330 3.330 3.330 23.33 6.670 6.670 6.670 16.67
Delimiters ASR(%) 73.33 80.00 80.00 53.33 90.00 96.67 93.33 40.00 76.67 76.67 80.00 33.33 6.670 10.00 20.00 60.67
Sandwich. ASR(%) 76.67 86.67 86.67 46.67 96.67 100.0 86.67 50.00 83.33 83.33 90.00 33.33 0.000 6.670 23.33 63.33
Instruction. ASR(%) 80.00 80.00 80.00 46.67 93.33 100.0 90.00 43.33 80.00 76.67 73.33 26.67 13.33 10.00 26.67 61.33
V-A2 Defense Performance
We evaluate 7 representative defense methods from both prevention-based and detection-based paradigms. These include perplexity filtering (PPL), windowed PPL (WinPPL), a naive LLM-based detector and retokenizationbaseDefense , delimiter insertion, sandwich prompting, and instruction augmentation. Since prevention-based and detection-based methods are based on different mechanisms and evaluated on other metrics, we report their results separately in Table VI and Table V, respectively.
Detection-based Defenses Among detection-based methods, perplexity-based filters (PPL and WinPPL) generally underperform. They achieve relatively low accuracy (48.35% and 51.22%), recall (53.11% and 62.44%), precision (44.98% and 48.66%), and F1 scores (0.484 and 0.545), indicating limited capability in distinguishing between adversarial and non-adversarial inputs. In contrast, the naive LLM-based Detector achieves the highest detection performance across all metrics, with an accuracy (65.67%), recall (64.71%), precision (65.49%), and F1 score (0.627), outperforming all other prevention-based defenses, particularly against complex heuristic-based manipulations. Its effectiveness stems from leveraging a powerful LLM to identify subtle adversarial patterns in responses. However, this approach incurs significant computational overhead, as it requires querying the LLM detector for each input response.
Prevention-based Defenses. Overall, prevention-based methods demonstrate moderate effectiveness across most attacks. However, they remain largely ineffective against stronger adversarial attacks, such as Combined Attack (H6) and AdvEval (H7). Among these methods, Retokenization yields the best average defense performance, with an ASR reduced to 16.67% and an iSDR of -0.363, significantly outperforming the other methods, which exhibit ASRs exceeding 60% and positive iSDR values above 0.03. Despite this relative success, prior findings baseDefense indicate that Retokenization can degrade model performance on benign inputs. This degradation arises from altered tokenization patterns, which can potentially affect how the model interprets the content and lead to inaccurate evaluations. Thus, while Retokenization offers strong adversarial robustness, it introduces a trade-off between defense effectiveness and benign input performance.